OpenStudy (anonymous):

Medals!!! Find the distance between the following points: (-3, 3)(-6, 7).

OpenStudy (anonymous):

you can use this \[D=\sqrt{(y _{b}-y _{a})^2+(x _{b}-x _{a})^2}\]

OpenStudy (anonymous):

which one is yb ya xb and xa?

OpenStudy (anonymous):

a=(-3,3) b=(-6,7) \[D=\sqrt{(-6+3)^2+(7-3)^2}\] \[\sqrt{9+16}=\sqrt{25}=5\] but if a=(-6,7) and b=(-3,3) the answer will be the same . \[D=\sqrt{(-3+6)^2+(3-7)^2}=\sqrt{9+16}=5\]

OpenStudy (anonymous):

so its 5?

OpenStudy (anonymous):

units

OpenStudy (anonymous):

@amirreza1870

OpenStudy (anonymous):

ummmm hello!!!!!

OpenStudy (anonymous):

most students usually say 5 cm. but I can say 5m and 5 km and ... so they usually don't write any thing after 5 . your Coordinate system can be very big .every square can be 1m^2.so distance between your points can be 5m.

OpenStudy (anonymous):

ok thank you!